# W-HMR: Human Mesh Recovery in World Space with Weak-supervised Camera Calibration and Orientation Correction

### v3

@@ 2.3 Orientation Correction
where $\oplus$ represents concatenation, $FC(\cdot)$ are couples of FC layers. To avoid distress, we explain why the operation of rotation matrices here does not use multiplication but rather direct addition. The network of OrientCorrect is similar to the classic iterative regressor used in HMR [14]. The similar design, instead of multiplication, has been employed in many previous works [25, 18, 54, 44] to predict rotations. The global information $glob\_info=\left\{\frac{\varDelta c_{x}}{\tilde{f}},\frac{\varDelta c_{y}}{\tilde{f}},\frac{h_{bbox}}{\tilde{f}},\frac{W}{\tilde{f}},\frac{H}{\tilde{f}}\,\,\right\}$ and it is important auxiliary information for predicting body orientation [25]. As for $\tilde{f}=\sqrt{W^{2}+H^{2}}$ in global information, it is just for normalization here and does not hold any geometric meanings, so we don’t use the predicted focal length here. The OrientCorrect is a completely separate module. So, it not only makes full use of what the model has learned in camera coordinate, but also avoids destroying the body feature space that the model has constructed. Additionally, this method can be easily applied to other models based on camera coordinate without reducing the original precision, which is shown in Table 8.
2.4 Training Paradigm
[…]
@@ 2.4 Training Paradigm
[…]
In the first two stages, training is conducted entirely in camera coordinate. In the third stage, we need to equip the model to reconstruct the human body in world coordinate. The pre-trained regressors, focal length predictor, and other modules are frozen at this stage. We only need to train the OrientCorrect, which is composed of three FC layers. We exclusively use the 3D joint, mesh vertices, and SMPL parameters in world coordinate for supervision.
2.5 Other Losses
@@
First, let’s delve into the IUV map, which is integral to the IUV loss. The IUV map, as defined in [10], uses "I" to indicate the body part to which a pixel belongs and "UV" to represent the 2D coordinates on the surface of the body part. This mapping establishes dense correspondences from 3D mesh to 2D plane. We employ three 2D convolution layers to predict the IUV map based on the feature map $\varPhi_{3}$ and render the ground truth online. The loss function is defined as
@@
3 Experiment
3.1 Implementation Details
Table 1: This table illustrates the changes in the size of input images and features when using different backbones (ResNet50 or VitPose-B) and employing different downsampling methods (sparse vertices or markers). As for what symbols mean, please refer to Section 3.1.
@@ 3.1 Implementation Details
Hybrid Regression Module For the spatial features $\{\phi_{i}^{s}\}_{i=1}^{3}$, we have corresponding regression models. We utilize parameter regressors for the first two spatial features to obtain the initial meshes. These regressors take spatial features $\phi_{i}^{s}$, global information $glob\_info$, and former SMPL parameters $\varTheta_{i-1}$ as inputs and output SMPL parameters $\varTheta_{i}$. The parameter regressors consist of several FC layers and iteratively produce the final result like
@@
Note $\varTheta_{0}$ is set as the mean $\overline{\varTheta}$ following [14]. As for the final regressor, we have two versions in total. The first version still involves the parameter regressor. The second version utilizes a vertex regressor to obtain a finer human body mesh. After obtaining a coarse human mesh using the first two regressors, we introduce a model-free vertex regressor based on the transformer. This vertex regressor takes spatial features $\phi_{3}^{s}$ and point-wise features $\left\{\phi_{3,n}^{p}\in\mathbb{R}^{C_{i}}\right\}_{n=1}^{N_{i}}$ as input. It’s important to note that no "$fc$" exists in this case, indicating that these features do not undergo channel reduction. To reduce network complexity and adapt to the SMPL model, instead of directly recovering the 6890 vertices defined by SMPL, we recover 431 sparse vertices and gradually restore them to 6890 vertices using two linear layers.
@@
We train three models in total for evaluation. Two of them utilize a hybrid regression module, using 431 sparse vertices for downsampling. The distinction of these two models lies in using different backbones, including ResNet50 [11] and VitPose-B [48]. Additionally, we train an extra model using a pure parameter regression module, employing 67 mocap markers for downsampling, with the backbone being VitPose-B. For training, we utilize datasets such as Human 3.6m [12], 3DPW [42], MPII3D [31], MPII [1], AGORA [35], COCO [28], SPEC-SYN [19], and HuMMan [3]. Some of the datasets are augmented with pseudo-labels generated by CamCalib, EFT [13], or CLIFF [25]. Please refer to Section 3.2 for detailed information about the datasets used. Our model can simultaneously provide results in camera and world coordinate. But if results in world coordinate are not needed or camera rotation is known, the CamCalib and OrientCorrect can be discarded. The focal length predictor can also be discarded if mesh projection is not required.
@@
We experiment with different backbones and downsampling points to validate our method and explore the optimal model structure. As a result, there are corresponding variations in the sizes of input images and features. We summarize these variations and create Table 1 to provide an overview of these changes. The training process is conducted with a batch size of 128 and a fixed learning rate of 0.00005 on three RTX 3090 GPUs until convergence, taking approximately one week. Following [54, 18], we apply standard data augmentation techniques such as rotation, flipping, scaling, translation, noise, and occlusion. As for hyperparameters, $\lambda_{1}=\lambda_{2}=\lambda_{3}=300$, $\lambda_{4}=\lambda_{5}=\lambda_{6}=15$, $\lambda_{7}=0.06$, $\lambda_{8}=60$, $\lambda_{9}=0.03$ and $\lambda_{10}=\lambda_{11}=\lambda_{12}=0.05$.
3.2 About Datasets
@@
In our experiments, we utilize the following datasets: Human 3.6M [12], 3DPW [42], MPII [1], COCO [28], MPI-INF-3D [31], AVA  [9], AIC [46], Insta [15], AGORA [35], SPEC-SYN [19], SPEC-MTP [19], and HuMMan [3]. We conduct experiments on the distorted datasets AGORA, HuMMan, and SPEC-MTP to evaluate our model. Furthermore, following traditional practices[45, 54], we also evaluate our model on 3DPW, Human 3.6M, and MPII-INF-3D. We employ AGORA’s evaluation platform as a standard for ablation experiments due to its diverse scenes, characters, and shooting methods. We believe that the evaluation results obtained on AGORA are more representative. For human recovery in world space, We additionally perform an ablation study on SPEC-SYN to validate the robustness of our method. Following are details about each dataset.
@@
3DPW’ [42] full name is "3D Human in the Wild." It is an outdoor dataset that captures actor information using motion capture devices. This dataset is known for its difficulty and has been widely used by researchers to demonstrate the robustness of their methods. Following past conventions, we train two models, one without 3DPW for training and another using 3DPW for training, to evaluate our model’s robustness.
@@
Human 3.6M [12] is widely used as a standard dataset to assess the estimation of 3D human pose and the recovery of human mesh. To reduce data redundancy, the original videos are down-sampled from 50 to 10 fps, resulting in 312,188 frames available for training and 26859 frames for evaluation. Following established protocols, we utilize five subjects (S1, S5, S6, S7, S8) for training and two subjects (S9, S11) for evaluation.
@@
MPII [1] is collected from YouTube and comprises a diverse range of activities in various scenes. Based on [53, 8], we only utilize the portion with pseudo 3D labels, which contains approximately 14,000 images.
@@
Figure 5: Qualitative results on AGORA compared with SPEC [19]. The areas marked within the red circles indicate the mismatch between the SPEC reconstructed mesh and the original image. In contrast, by comparing it with the corresponding W-HMR reconstructed results enclosed in the yellow circles, it is evident that our approach exhibits stronger capabilities in handling complex samples.
Table 2: Comparison with SOTA methods on distorted datasets in camera coordinate. Some of the evaluation results on HuMMan [3] and SPEC-MTP [19] are sourced from  [43], while the evaluation results on AGORA [35] are obtained from its official website https://agora-evaluation.is.tuebingen.mpg.de/. Including the following tables, $*{}^{*}$ represents backbone is ResNet50 [11], ${}^{\dagger}$ means VitPose-B [48, 49], and "P" denotes the pure parameter module. Note we do not perform any fine-tuning on SPEC-MTP.
@@
$*{}^{*}$ 78.1 87.0 70.3 78.3 37.4 29.1 40.5 104.5 61.7 118.7
${}^{\dagger}$ 68.7 77.7 61.8 69.9 39.2 30.8 40.4 104.9 62.3 115.2
${}^{\dagger}$ 70.4 75.4 63.4 67.9 36.5 27.1 43.5 108.1 65.7 121.5
@@
AGORA [35] is a multi-person synthetic dataset that has gained recognition in recent years due to its precise SMPL-X labels, complex characters and scenes, and fair evaluation methods. The dataset is available in two resolutions: 1280x720 and 3840x2160. We exclusively utilize the low-resolution set in our work. The train, validation, and test sets are used according to their original partitioning. The bounding boxes used in the test set are obtained using the 2D keypoints generated by Hand4Whole [33].
@@ 3.3 About Metrics
As for the metrics in world coordinate, we follow SPEC [19] using W-MPJPE, PA-MPJPE, and W-PVE. PA-MPJPE (Procrustes-aligned Mean Per Joint Position Error) exists as a metric that purely evaluates the accuracy of body reconstruction. It removes the effect of unknown camera rotation by Procrustes-aligned and has therefore been adopted by many previous works. The calculation of W-MPJPE and W-PVE is the same as MPJPE and PVE. The only difference is that the labels used are adjusted to the ground truth in world coordinate. We, including SPEC, believe these metrics represent the model’s capabilities in real-world applications.
3.4 Evaluation Results
@@
Camera Coordinate In our study, we choose three challenging datasets containing distorted images to showcase the superiority of weak-supervised camera calibration under complex conditions. As depicted in Table 2, our method surpasses the performance of most existing methods on the challenging synthetic dataset AGORA [35]. Each of our models demonstrates unique strengths. The pure regression model achieves higher accuracy in joint localization, and the hybrid regression module excels in capturing mesh details. Compared to baseline [54], our improvement is quite substantial. Note the baseline’s backbone, training datasets and evaluation settings are the same as $*{}^{*}$. As depicted in Fig. 5, our W-HMR outperforms SPEC [19] in handling occlusions, rare body rotations, and body intersections. Although the joint accuracy does not exhibit significant improvement on the indoor dataset HuMMan [3], we achieve the closest performance compared to the latest method Zolly in PVE, even surpassing the model-free method FastMETRO [4]. The most remarkable performance is observed on the real distorted dataset SPEC-MTP. We significantly elevate the SOTA results on SPEC-MTP to new heights.
Table 3: Comparison with SOTA methods on SPEC-MTP [19]. The first six models in the table are all methods that output results in camera coordinate, so there are two results. The "a" in "a/b" indicates the result calculated directly using the output in the camera coordinate. The "b" is the result calculated after the output has been calibrated by camera rotation predicted by CamCalib. The latter two models can directly get the results in world coordinate without calibration, so there is only one result.
@@
${}^{\dagger}$ 118.7 66.6 133.9
World Coordinate We utilize SPEC-MTP [19] as the primary evaluation dataset. Due to its only test set, specific fine-tuning is not feasible. It adopts a shooting method different from traditional datasets, resulting in unique rotation and severe distortion. Therefore, it poses a significant challenge and is particularly suitable for demonstrating OrientCorrect’s strength. To ensure fairness, we only evaluate the W-HMR with a pure parameter regression module. W-HMR easily outperforms existing methods, including SPEC. It is worth noting that the superior performance of W-HMR in world coordinates is not entirely dependent on higher reconstruction accuracy. Perhaps you noticed that there are two Ours-Ps in Table 3. This is because W-HMR can generate results in two coordinates simultaneously. By comparing these two sets of results, we can conclude that OrientCorrect is significantly superior to the traditional method of directly using predicted camera rotation for correction. Upon comparing our results calibrated by CamCalib with those corrected by OrientCorrect, it is evident that our module is effectively functioning. As illustrated in W-HMR: Human Mesh Recovery in World Space with Weak-supervised Camera Calibration and Orientation Correction, the reconstructed human body pose appears implausible in world space when the camera coordinate is not aligned with the world coordinate. However, our OrientCorrect corrects the orientation of the human body and maintains its reasonableness in world coordinate. Our model can accurately perceive and correct both slight and significant camera rotation. Quantitative and qualitative results both demonstrate that our method is simpler and more efficient than SPEC, which will be further substantiated in subsequent ablation studies.
Table 4: Comparison with SOTA Methods on 3DPW [42]. The column "w/ 3DPW" indicates whether the model has been trained using the 3DPW dataset. We bold both the best results with and without 3DPW.
@@
$*{}^{*}$ No 75.6 46.7 87.7
${}^{\dagger}$ No 85.4 54.6 96.0
${}^{\dagger}$ No 82.9 52.1 95.1
$*{}^{*}$ Yes 68.5 43.2 81.6
${}^{\dagger}$ Yes 64.6 40.5 75.7
${}^{\dagger}$ Yes 67.3 41.6 79.9
Table 5: Comparison with SOTA Methods on controllable datasets including Human 3.6M [12] and MPI-INF-3D [31].
@@
Traditional Benchmarks Following previous methods [54, 45], we conduct evaluation on 3DPW [42], H36M [12], and MPII-INF-3D [31]. Specifically, we evaluate the performance of models trained with and without the 3DPW train set on the 3DPW test set. As shown in Table 4, we achieve remarkable improvements on this challenging in-the-wild dataset, surpassing previous works easily. Even compared to the SOTA method Zolly, W-HMR still exhibits significant advancements when trained without the 3DPW train set, which shows W-HMR’s robustness. In Table 5, W-HMR’s performance is relatively less impressive in the case of controllable datasets, although it remains competitive with these SOTA methods. Controllable datasets refer to datasets obtained under laboratory conditions, typically characterized by limited variations in actors and scenes, which are prone to overfitting. HuMMan in Table 2 is also an example of such datasets. So, it becomes difficult for our advantages to be fully demonstrated.
@@
Qualitative Results More visualizations are generated to give a more comprehensive reference. Also, we welcome running our codes on your samples. Back to qualitative results, from Fig. 7, we can see that our model is able to recover the human body very well when facing various scenes with various characters in multiple poses. We purposely show the results from side angles and in world space. In the past, many methods could also make the reconstructed mesh match the original image perfectly, but it would not look very good if we changed the visual angle. As shown in Fig. 7, W-HRM doesn’t suffer from this issue and performs well in every condition.
Figure 6: Qualitative results on HuMMan [3]. "Input" refers to input images of the model. "Overlay" means we project the recovered mesh back onto the original images. "Cam" and "World" show the output in camera and world coordinates, respectively.
@@
Figure 7: Qualitative results in camera and world coordinates. W-HMR shows excellent reconstruction accuracy and robustness. It is able to reconstruct various human bodies well in different scenes. Not only does the reconstructed mesh align well with the original image, but even when switching the angle of view, the reconstructed human body can still be seen to fit the target human.
@@
Methods based on camera coordinate often fail to meet the needs of applications in world space. It may not be noticeable from Fig. 7 because they fit the weak-perspective camera model and the assumption of $R_{c}=I$. But from Fig. 6, the drawbacks of the traditional method are undeniable. The reconstructed human body is entirely lopsided in world coordinate. Instead, our OrientCorrect module can correct the pose well to achieve a reasonable pose recovery in world coordinate. This method has a wide range of applications, allowing the model to reconstruct a proper human body when facing in-the-wild images. Before W-HMR, there were also methods like SPEC that explored human mesh recovery in world space. As a pioneering work, SPEC opened up a new way of thinking for subsequent research, but the performance was not robust enough. In Fig. 8, we can see that SPEC is not as accurate as W-HMR in any coordinate, showing that our method cleverly avoids the error accumulation of SPEC and achieves all-around transcendence.
@@
Figure 8: Qualitative comparison with SPEC on SPEC-MTP [19]. Here are some examples: W-HMR shows better mesh-image alignment and higher orientation accuracy in world space.
3.5 Ablation Study

### v4

@@ III-C Orientation Correction
where $\oplus$ represents concatenation, $FC(\cdot)$ are only couples of FC layers, $R_{c}$ is the camera rotation output by CamClib, $R_{b}^{c}$ is the predicted body orientation $\theta_{0}$, and $R_{b}$ is the final predicted body orientation in the world coordinate. Under the supervision of true body orientation in the world coordinate, $FC(\cdot)$ learns to correct the $R_{b}^{c}$ to $R_{b}$. No additional operation is required to ensure the output rotation matrix $R_{b}$ is valid. To avoid distress, we explain why the operation of rotation matrices here does not use multiplication but rather direct addition. The network of OrientCorrect is similar to the classic iterative regressor used in HMR [1]. A similar design, instead of multiplication, has been employed in many previous works [4, 11, 13, 34] to predict rotations. The global information $glob\_info=\left\{\frac{\varDelta c_{x}}{\tilde{f}},\frac{\varDelta c_{y}}{\tilde{f}},\frac{h_{bbox}}{\tilde{f}},\frac{W}{\tilde{f}},\frac{H}{\tilde{f}}\,\,\right\}$ and it is important auxiliary information for predicting body orientation [4]. As for $\tilde{f}=\sqrt{W^{2}+H^{2}}$ in global information, it is just for normalization here and does not hold any geometric meanings, so we do not use the predicted focal length here. The OrientCorrect is a completely separate module. So, it not only makes full use of what the model has learned in the camera coordinate but also avoids destroying the body feature space that the model has constructed. Additionally, this method can be easily applied to other models based on the camera coordinate without reducing the original precision, shown in Tab. VIII.
III-D Training Paradigm
As shown in Fig. 5, our training paradigm is divided into three stages. In stage I, we mainly train the regressors and focal predictors of the model to give them initial predictive capabilities. In stage II, we use FULL2 2D supervision to improve accuracy further. In stage III, we train OrientCorrect alone to give W-HMR the ability to recover humans in world space. The following is a detailed description of each stage.
In the first stage, after obtaining the 3D joints, we need to project them onto both the cropped image and the original image to obtain $J_{2D}$ and $J_{2D}^{full}$, respectively. $J_{2D}$ still adheres to the traditional weak-perspective camera model, while $J_{2D}^{full}$ uses the full-perspective camera model. During the FULL2 2D supervision, we detach the gradients of the 3D joints, scale $s$, and translation $(t_{x},\,t_{y})$ to prevent incorrect backpropagation. This precaution is necessary because, in the first stage, the focal length predictor cannot predict a reasonable focal length. This would lead to a completely erroneous 2D joint loss, affecting the regressors to learn to reconstruct the human body. Moreover, to stabilize the training of the focal length predictor, we introduce an additional loss in the first stage, i.e.,
@@ III-D Training Paradigm
Since the model has already developed initial capabilities in predicting joints and focal lengths in the first stage, we can directly apply FULL2 2D supervision without concern about training collapse. We continue to detach $(s,\,t_{x},\,t_{y})$ from the full-perspective projection. Because the perspective effect does not impact the model’s learning of scale and translation. In the second stage, we detach the gradients of the 3D joints from the weak-perspective projection to prevent erroneous backpropagation of the weak-perspective 2D joint loss on cropped images.
Training is conducted entirely in the camera coordinate in the first two stages. In the third stage, we need to equip the model to reconstruct the human body in the world coordinate. The pre-trained regressors, focal length predictor, and other modules are frozen. We only need to train the OrientCorrect, composed of three FC layers. We exclusively use the 3D joint, mesh vertices, and SMPL parameters in the world coordinate for supervision.
III-E Other Losses
@@
First, let’s delve into the IUV map, which is integral to the IUV loss. The IUV map, as defined in [35], uses ”I” to indicate the body part to which a pixel belongs and ”UV” to represent the 2D coordinates on the surface of the body part. This mapping establishes dense correspondences from 3D mesh to 2D plane. We employ three 2D convolution layers to predict the IUV map based on the feature map $\varPhi_{3}$ and render the ground truth online. The loss function is defined as
@@
TABLE I: This table illustrates the changes in the size of input images and features when using different backbones (ResNet50 or VitPose-B) and employing different downsampling methods (sparse vertices or markers). As for what symbols mean, please refer to Sec. IV-A.
@@ IV-A Implementation Details
Hybrid Regression Module For the spatial features $\{\phi_{i}^{s}\}_{i=1}^{3}$, we have corresponding regression models. We utilize parameter regressors for the first two spatial features to obtain the initial motion. These regressors take spatial features $\phi_{i}^{s}$, global information $glob\_info$, and former SMPL parameters $\varTheta_{i-1}$ as inputs and output SMPL parameters $\varTheta_{i}$. The parameter regressors consist of several FC layers and iteratively produce the final result like
@@
Note $\varTheta_{0}$ is set as the mean $\overline{\varTheta}$ following [1]. As for the final regressor, we have two versions in total. The first version still involves the parameter regressor. The second version utilizes a vertex regressor to obtain a finer human body motion. After obtaining a coarse human motion using the first two regressors, we introduce a model-free vertex regressor based on the transformer. This vertex regressor takes spatial features $\phi_{3}^{s}$ and point-wise features $\left\{\phi_{3,n}^{p}\in\mathbb{R}^{C_{i}}\right\}_{n=1}^{N_{i}}$ as input. It’s important to note that no ”$fc$” exists in this case, indicating that these features do not undergo channel reduction. To reduce network complexity and adapt to the SMPL model, instead of directly recovering the 6890 vertices defined by SMPL, we recover 431 sparse vertices and gradually restore them to 6890 vertices using two linear layers.
@@
We train three models in total for evaluation. Two of them utilize a hybrid regression module, using 431 sparse vertices for downsampling. The distinction between these two models lies in using different backbones, including ResNet50 [36] and VitPose-B [37]. Additionally, we train an extra model using a pure parameter regression module, employing 67 mocap markers for downsampling, with the backbone being VitPose-B. For training, we utilize datasets such as Human 3.6m [38], 3DPW [39], MPII3D [40], MPII [41], AGORA [42], COCO [43], SPEC-SYN [3], and HuMMan [44]. Some datasets are augmented with pseudo-labels generated by CamCalib, EFT [45], or CLIFF [4]. Please refer to Sec. IV-B for detailed information about the datasets used. Our model can simultaneously provide results in camera and world coordinates. But if results in the world coordinate are not needed or camera rotation is known, the CamCalib and OrientCorrect can be discarded. The focal length predictor can also be discarded if mesh projection is not required.
@@
We experiment with different backbones and downsampling points to validate our method and explore the optimal model structure. As a result, there are corresponding variations in the sizes of input images and features. We summarize these variations and create Tab. I to provide an overview of these changes. The training process is conducted with a batch size of 128 and a fixed learning rate of 0.00005 on three RTX 3090 GPUs until convergence, taking approximately one week. Following [13, 11], we apply standard data augmentation techniques such as rotation, flipping, scaling, translation, noise, and occlusion. As for hyperparameters, $\lambda_{1}=\lambda_{2}=\lambda_{3}=300$, $\lambda_{4}=\lambda_{5}=\lambda_{6}=15$, $\lambda_{7}=0.06$, $\lambda_{8}=60$, $\lambda_{9}=0.03$ and $\lambda_{10}=\lambda_{11}=\lambda_{12}=0.05$.
IV-B About Datasets
@@
In our experiments, we utilize the following datasets: Human 3.6M [38], 3DPW [39], MPII [41], COCO [43], MPI-INF-3D [40], AVA  [46], AIC [47], Insta [48], AGORA [42], SPEC-SYN [3], SPEC-MTP [3], and HuMMan [44]. To evaluate our model, we conduct experiments on the distorted datasets AGORA, HuMMan, and SPEC-MTP. Furthermore, following traditional practices[49, 13], we also evaluate our model on 3DPW, Human 3.6M, and MPII-INF-3D. We employ AGORA’s evaluation platform as a standard for ablation experiments due to its diverse scenes, characters, and shooting methods. We believe that the evaluation results obtained on AGORA are more representative. For human recovery in world space, We additionally perform an ablation study on SPEC-SYN to validate the robustness of our method. Following are details about each dataset.
@@
3DPW’ [39] full name is ”3D Human in the Wild.” It is an outdoor dataset that captures actor information using motion capture devices. This dataset is known for its difficulty and has been widely used by researchers to demonstrate the robustness of their methods. Following past conventions, we train two models, one without 3DPW for training and another using 3DPW for training, to evaluate our model’s robustness.
@@
Human 3.6M [38] is widely used as a standard dataset to assess the estimation of 3D human pose and the recovery of human motion. To reduce data redundancy, the original videos are down-sampled from 50 to 10 fps, resulting in 312,188 frames available for training and 26859 frames for evaluation. Following established protocols, we utilize five subjects (S1, S5, S6, S7, S8) for training and two subjects (S9, S11) for evaluation.
@@
MPII [41] is collected from YouTube and comprises various activities in various scenes. Based on [7, 14], we only utilize the portion with pseudo 3D labels, containing approximately 14,000 images.
@@
TABLE II: Comparison with SOTA methods on distorted datasets in the camera coordinate. Some of the evaluation results on HuMMan [44] and SPEC-MTP [3] are sourced from  [2], while the evaluation results on AGORA [42] are obtained from its official website https://agora-evaluation.is.tuebingen.mpg.de/. Including the following tables, represents backbone is ResNet50 [36], means VitPose-B [37, 52], and ”P” denotes the pure parameter module. Note we do not perform any fine-tuning on SPEC-MTP.
@@
Ours 78.1 87.0 70.3 78.3 37.4 29.1 40.5 104.5 61.7 118.7
Ours 68.7 77.7 61.8 69.9 39.2 30.8 40.4 104.9 62.3 115.2
Ours-P 70.4 75.4 63.4 67.9 36.5 27.1 43.5 108.1 65.7 121.5
Figure 6: Qualitative results on AGORA compared with SPEC [3]. The areas marked within the red circles indicate the mismatch between the SPEC reconstructed motion and the original image. In contrast, by comparing it with the corresponding W-HMR reconstructed results enclosed in the yellow circles, it is evident that our approach exhibits stronger capabilities in handling complex samples.
@@
AGORA [42] is a multi-person synthetic dataset that has gained recognition recently due to its precise SMPL-X labels, complex characters and scenes, and fair evaluation methods. The dataset is available in two resolutions: 1280x720 and 3840x2160. We exclusively utilize the low-resolution set in our work. The train, validation, and test sets are used according to their original partitioning. The bounding boxes used in the test set are obtained using the 2D keypoints generated by Hand4Whole [57].
@@ IV-C About Metrics
As for the metrics in the world coordinate, we follow SPEC [3] using W-MPJPE, PA-MPJPE, and W-PVE. PA-MPJPE (Procrustes-aligned Mean Per Joint Position Error) exists as a metric that purely evaluates the accuracy of body reconstruction. It removes the effect of unknown camera rotation by Procrustes-aligned and has therefore been adopted by many previous works. The calculation of W-MPJPE and W-PVE is the same as that of MPJPE and PVE. The only difference is that the labels used are adjusted to the ground truth in the world coordinate. We, including SPEC, believe these metrics represent the model’s capabilities in real-world applications.
IV-D Evaluation Results
@@
Camera Coordinate In our study, we choose three challenging datasets containing distorted images to showcase the superiority of weak-supervised camera calibration under complex conditions. As depicted in Tab. II, our method surpasses the performance of most existing methods on the challenging synthetic dataset AGORA [42]. Each of our models demonstrates unique strengths. The pure regression model achieves higher accuracy in joint localization, and the hybrid regression module excels in capturing mesh details. Compared to baseline [13], our improvement is quite substantial. Note the baseline’s backbone, training datasets, and evaluation settings are the same as Ours. As depicted in Fig. 6, our W-HMR outperforms SPEC [3] in handling occlusions, rare body rotations, and body intersections. Although the joint accuracy does not exhibit significant improvement on the indoor dataset HuMMan [44], we achieve the closest performance compared to the latest method Zolly in PVE, even surpassing the model-free method FastMETRO [54]. The most remarkable performance is observed on the real distorted dataset SPEC-MTP. We significantly elevate the SOTA results on SPEC-MTP to new heights.
TABLE III: Comparison with SOTA methods on SPEC-MTP [3]. The first six models in the table are all methods that output results in the camera coordinate, so there are two results. The ”a” in ”a/b” indicates the result calculated directly using the output in the camera coordinate. The ”b” is the result calculated after camera rotation predicted by CamCalib has calibrated the output. The latter two models can directly get the results in the world coordinate without calibration, so there is only one result.
@@
Ours-P 118.7 66.6 133.9
World Coordinate We utilize SPEC-MTP [3] as the primary evaluation dataset. Due to its only test set, specific fine-tuning is not feasible. It adopts a shooting method different from traditional datasets, resulting in unique rotation and severe distortion. Therefore, it poses a significant challenge and is suitable for demonstrating OrientCorrect’s strength. We only evaluate the W-HMR with a pure parameter regression module to ensure fairness. W-HMR easily outperforms existing methods, including SPEC. It is worth noting that the superior performance of W-HMR in the world coordinates is not entirely dependent on higher reconstruction accuracy. Perhaps you noticed two Ours-Ps in Tab. III. This is because W-HMR can generate results in two coordinates simultaneously. By comparing these two sets of results, we can conclude that OrientCorrect is significantly superior to the traditional method of directly using predicted camera rotation for correction. Upon comparing our results calibrated by CamCalib with those corrected by OrientCorrect, it is evident that our module is effectively functioning. As illustrated in Fig. 1, the reconstructed human body pose appears implausible in world space when the camera coordinate is not aligned with the world coordinate. However, our OrientCorrect corrects the orientation of the human body and maintains its reasonableness in the world coordinate. Our model can accurately perceive and correct both slight and significant camera rotation. Quantitative and qualitative results demonstrate that our method is simpler and more efficient than SPEC, which will be further substantiated in subsequent ablation studies.
Figure 7: Qualitative results on HuMMan [44]. ”Input” refers to input images of the model. ”Overlay” means we project the recovered mesh back onto the original images. ”Cam” and ”World” show the output in camera and world coordinates, respectively.
TABLE IV: Comparison with SOTA Methods on 3DPW [39]. The column ”w/ 3DPW” indicates whether the model has been trained using the 3DPW dataset. We bold both the best results with and without 3DPW.
@@
Ours No 75.6 46.7 87.7
Ours No 85.4 54.6 96.0
Ours-P No 82.9 52.1 95.1
Ours Yes 68.5 43.2 81.6
Ours Yes 64.6 40.5 75.7
Ours-P Yes 67.3 41.6 79.9
TABLE V: Comparison with SOTA Methods on controllable datasets including Human 3.6M [38] and MPI-INF-3D [40].
@@
Traditional Benchmarks Following previous methods [13, 49], we conduct evaluation on 3DPW [39], H36M [38], and MPII-INF-3D [40]. Specifically, we evaluate the performance of models trained with and without the 3DPW train set on the 3DPW test set. As shown in Tab. IV, we achieve remarkable improvements on this challenging in-the-wild dataset, surpassing previous works easily. Even compared to the SOTA method Zolly, W-HMR still exhibits significant advancements when trained without the 3DPW train set, which shows W-HMR’s robustness. In Tab. V, W-HMR’s performance is relatively less impressive in the case of controllable datasets, although it remains competitive with these SOTA methods. Controllable datasets refer to datasets obtained under laboratory conditions, typically characterized by limited variations in actors and scenes, which are prone to overfitting. HuMMan in Tab. II is an example of such datasets. So, it becomes difficult for our advantages to be fully demonstrated.
@@
Qualitative Results More visualizations are generated to give a more comprehensive reference. Also, we welcome running our codes on your samples. Back to qualitative results, from Fig. 8, we can see that our model can recover the human body very well when facing various scenes with various characters in multiple poses. We purposely show the results from side angles and in world space. In the past, many methods could also make the reconstructed mesh match the original image perfectly, but it would not look perfect if we changed the visual angle. As shown in Fig. 8, W-HRM does not suffer from this issue and performs well in every condition.
@@
Figure 8: Qualitative results in camera and world coordinates. W-HMR shows excellent reconstruction accuracy and robustness. It can reconstruct various human bodies well in different scenes. Not only does the reconstructed mesh align well with the original image, but even when switching the angle of view, the reconstructed human body can still be seen to fit the target human.
@@
Methods based on the camera coordinate often fail to meet the needs of applications in world space. It may not be noticeable from Fig. 8 because they fit the weak-perspective camera model and the assumption of $R_{c}=I$. But from Fig. 7, the drawbacks of the traditional method are undeniable. The reconstructed human body is entirely lopsided in the world coordinate. Instead, our OrientCorrect module can correct the pose well to achieve a reasonable pose recovery in the world coordinate. This method has many applications, allowing the model to reconstruct a proper human body when facing in-the-wild images. Before W-HMR, there were also methods like SPEC that explored human motion recovery in world space. As a pioneering work, SPEC opened up a new way of thinking for subsequent research, but the performance was not robust enough. In Fig. 9, we can see that SPEC is not as accurate as W-HMR in any coordinate, showing that our method cleverly avoids the error accumulation of SPEC and achieves all-around transcendence.
@@
Figure 9: Qualitative comparison with SPEC on SPEC-MTP [3]. Here are some challenging examples. Although imperfect, W-HMR shows better mesh-image alignment and higher orientation accuracy in world space.
Camera Parameter There is no such evaluation for inferred camera parameters because W-HMR cannot and does not need to predict true intrinsic and extrinsic parameters. First, focal length prediction is an ill-posed problem and lacks much training data. We choose to leverage mesh-image alignment information to predict pseudo focal lengths. This pseudo focal length may differ from the true focal length but can accurately project the recovered mesh to the human area. After all, the purpose of predicting the focal length is to handle distorted images so that the calculation of 2D joint loss is more accurate, thereby improving model performance. In summary, inferred camera parameters cannot be evaluated quantitatively because they may differ significantly from ground truth. Of course, it can be implicated by improved recovery accuracy and aligned-well qualitative results. Moreover, the module used to predict the camera rotation matrix in W-HMR comes from CamClib in SPEC [3]. Its accuracy has been evaluated in SPEC. The novelty in our work is how we better use this information. With the same CamClib, W-HMR still significantly outperforms SPEC, which is a solid evidence to prove the effectiveness of our method.
IV-E Ablation Study